# k* Distribution: Evaluating the Latent Space of Deep Neural Networks using Local Neighborhood Analysis

### v1

@@ IV Experimental Results and Visualization
IV-A Common Experimental Setup
Datasets: Initially, we assess the k* distribution using the entire set of $1280$ images from the $16$-class-ImageNet dataset curated by Geirhos et al. [44]. This dataset contains $80$ images per each of the $16$ classes of the dataset. These are the $16$ entry-level categories from MS-COCO that have the highest number of ImageNet classes mapped via the WordNet hierarchy making them compatible with the $1,000$ classes of ImageNet-1k dataset [45]. The purpose of employing this dataset is to evaluate individual classes organized naturally.
In addition to the $16$-class-ImageNet dataset, we evaluate k* distribution comprehensively using $50,000$ validation images of the original ImageNet-1k dataset [45]. This dataset is divided into $50$ images for each of its $1,000$ classes of the dataset.
To further scrutinize the models’ latent spaces, we apply various transformations to the samples from both the $16$-class-ImageNet dataset and ImageNet-1k dataset. These transformations include Image Crop, the addition of Gaussian Noise, Rotation of the images, and the generation of Adversarial and Stylized Versions of the samples.
Moreover, to assess the k* distribution across various other tasks, we also use various other datasets for different tasks. For intent classification, we use the English subset of the MASSIVE [46] dataset. The testing samples contained $2,970$ samples split across $60$ distinct intent classes. For the keyword spotting task, we used the Speech Commands v0.02 dataset [47], which contains $4,890$ samples split across $36$ command categories. We utilize these datasets to analyze as described in Figure 2 and comprehend the characteristics associated with each pattern outlined in Figure 1.
Deep Neural Network Architectures: We analyze the latent space of pre-trained weights from various open-source neural networks referenced as we have used them in the experiments. Adversarial Attacks: We subject the trained ResNet-50 model to a Projected Gradient Descent (PGD) attack, as outlined in [48], employing a perturbation magnitude of ${\epsilon}=4/255$. This attack is aimed at gauging the extent of the effective alteration in representation induced by the adversarial samples.
Furthermore, we assess the model that has undergone adversarial training using the PGD attack. This evaluation aims to compare and contrast the latent space of the robust model with that of its non-robust counterpart.
TABLE I: Multi-category evaluation of various neural architectures by using statistical metrics across object categories. Metrics include Mean ($\mu_{k^{*}}$), Standard deviation ($\sigma_{k^{*}}$), Skewness Coefficient ($\gamma_{k^{*}}$), Accuracy (Acc), and prevailing Pattern (Pat) in the context of diverse visual patterns.
@@ IV-A Common Experimental Setup
TABLE II: Performance by various neural architectures across varied visual patterns - Pattern A (🟊normal-$\bigstar$), Pattern B (normal-$\clubsuit$), Pattern C (normal-$\spadesuit$), and $1,000$ ImageNet1k Classes. The table presents key statistics, including an average of Mean ($\mu_{k^{*}}$), Skewness coefficient ($\gamma_{k^{*}}$), Accuracy (Acc), and Number of Classes (N).
Average of Classes with Pattern A (🟊normal-$\bigstar$) (Fractured) Average of Classes with Pattern B (normal-$\clubsuit$) (Overlapped) Average of Classes with Pattern C (normal-$\spadesuit$) (Clustered) Average of 1,000 ImageNet1k Classes
@@
TABLE III: Multi-category evaluation of various layers of ResNet-50 by using statistical metrics across object categories. Metrics include Mean ($\mu_{k^{*}}$), Standard deviation ($\sigma_{k^{*}}$), Skewness Coefficient ($\gamma_{k^{*}}$), and prevailing Pattern (Pat) in the context of diverse visual patterns.
@@
TABLE IV: Performance by different layers of ResNet-50 across varied visual patterns - Pattern A (🟊normal-$\bigstar$), Pattern B (normal-$\clubsuit$), Pattern C (normal-$\spadesuit$), and $1,000$ ImageNet1k Classes. The table presents key statistics, including an average of Mean ($\mu_{k^{*}}$), Skewness coefficient ($\gamma_{k^{*}}$), and Number of Classes (N).
Average of Classes with Pattern A (🟊normal-$\bigstar$)  (Fractured) Average of Classes with Pattern B (normal-$\clubsuit$)  (Overlapped) Average of Classes with Pattern C (normal-$\spadesuit$)  (Clustered) Average of 1,000 ImageNet1k Classes
@@
TABLE V: Multi-category evaluation of models trained on different training distributions by using statistical metrics across object categories. Metrics include Mean ($\mu_{k^{*}}$), Standard deviation ($\sigma_{k^{*}}$), Skewness Coefficient ($\gamma_{k^{*}}$), Accuracy (Acc), and prevailing Pattern (Pat) in the context of diverse visual patterns.
@@
TABLE VI: Performance by models trained on different training datasets across varied visual patterns - Pattern A (🟊normal-$\bigstar$), Pattern B (normal-$\clubsuit$), Pattern C (normal-$\spadesuit$), and $1,000$ ImageNet Classes. The table presents key statistics, including an average of Mean ($\mu_{k^{*}}$), Skewness coefficient ($\gamma_{k^{*}}$), Accuracy (Acc), and Number of Classes (N).
Average of Classes with Pattern A (🟊normal-$\bigstar$) (Fractured) Average of Classes with Pattern B (normal-$\clubsuit$) (Overlapped) Average of Classes with Pattern C (normal-$\spadesuit$) (Clustered) Average of 1,000 ImageNet1k Classes
@@
TABLE VII: Multi-category evaluation of Robust (Adversarially Trained: AT) and Non-Robust (Standard Trained: ST) architectures by using statistical metrics across object categories. Metrics include Mean ($\mu_{k^{*}}$), Standard deviation ($\sigma_{k^{*}}$), Skewness Coefficient ($\gamma_{k^{*}}$), and prevailing Pattern (Pat) in the context of diverse visual patterns.
@@
TABLE VIII: Performance of Robust (Adversarially Trained: AT) and Non-Robust (Standard Trained: ST) architectures across varied visual patterns - Pattern A (🟊normal-$\bigstar$), Pattern B (normal-$\clubsuit$), Pattern C (normal-$\spadesuit$), and $1,000$ ImageNet1k Classes. The table presents key statistics, including an average of Mean ($\mu_{k^{*}}$), Skewness coefficient ($\gamma_{k^{*}}$), Accuracy (Acc), and Number of Classes (N).
Average of Classes with Pattern A (🟊normal-$\bigstar$) (Fractured) Average of Classes with Pattern B (normal-$\clubsuit$) (Overlapped) Average of Classes with Pattern C (normal-$\spadesuit$) (Clustered) Average of 1,000 ImageNet1k Classes
@@
IV-B Analysis of Latent Space of Different Neural Architectures
We have the option to visualize the latent space of ResNet-50 in two ways: using k* distribution, as illustrated in Figure 4 (Left), or employing dimensionality reduction techniques, showcased in Figure 4 (Right). Intriguingly, the t-SNE (Top Left) and UMAP (Bottom Right) visualizations indicate a highly clustered distribution of samples in latent space for ResNet-50, while Isomap (Top Right) and PCA (Bottom Left) show an overlapping distribution of samples in latent space. This creates uncertainty in distinguishing which classes are well-represented and which ones are fragmented in the distribution of samples in latent space.
To address this ambiguity and gain a clearer understanding of the local latent space, we turn to the k* distribution, visualized in Figure 4. By utilizing the k* distribution, we can distinctly identify that the local spaces of the six visualized classes differ. For instance, the Airplane and Elephant classes exhibit more Pattern C (normal-$\spadesuit$) clustered distribution of samples in latent space with a negatively skewed k* distribution.
In contrast, Bicycle and Dog classes showcase Pattern A (🟊normal-$\bigstar$) fractured distribution of samples in latent space with a positively skewed k* distribution, while Car and Truck have Pattern B (normal-$\clubsuit$) overlapped distribution of samples in latent space with an almost uniform k* distribution.
Furthermore, we can perform a comprehensive comparison of various neural architectures, as presented in Table I. This allows us to observe how specific classes are distributed differently across architectures. Notably, the Bicycle class displays Pattern A (🟊normal-$\bigstar$) (fractured distribution of samples) in ResNet and ResNeXt architectures, Pattern B (normal-$\clubsuit$)  (overlapped distribution of samples) in ViT, and Pattern C (normal-$\spadesuit$) (clustered distribution of samples) in EfficientNet-B0. Similarly, the Truck class is Pattern B (normal-$\clubsuit$) (overlapped distribution of samples) in ResNet and ResNeXt architectures, and Pattern C (normal-$\spadesuit$) (clustered distribution of samples) in ViT and EfficientNet-B0. Conversely, Airplane, Dog, and Elephant classes exhibit consistent distribution across all architectures.
Additionally, a comprehensive comparison of neural architectures can be made by calculating the averages of Fractured, Overlapped, and Clustered classes, as detailed in Table II. The results indicate distinct distributions of classes across various neural architectures. For instance, ResNet tends to fracture the latent space, with only 9 clustered classes. In contrast, ViT leans towards clustering of samples in the latent space, with a significant number of classes ($243$) exhibiting clustering.
A general trend is also observed, that there is an increase in the mean of k* distribution $\mu_{k^{*}}$ with improvements in model accuracy. Additionally, a declining trend in the skewness coefficient of k* distribution $\gamma_{k^{*}}$ suggests a transition from fractured to more overlapped distribution of samples in latent space across models.
IV-C Analysis of Latent Space of Different Layers of a Network
To gain insights into the representation of classes across different layers of the same model, we conduct an analysis of the latent space within various layers of ResNet-50. Specifically, we examine the latent space of the final logit layer, the average pooling layer, and several convolution layers situated at the conclusion of different stages in the ResNet-50 architecture, as depicted in Tables III and IV.
Upon observation, we note that the final logit layer induces changes in the representation of certain overlapped classes, transforming them from an average pooling state to predominantly clustered and often fractured representations. Additionally, we note that layers proximal to the input exhibit a highly fractured distribution of samples in latent space. This suggests that layers closer to the input aim to identify abstract patterns in the data, irrespective of the sample’s label as also suggested by Hohman et al. [24]. We also note the association of samples to the supervised class label increases through the network with the best association occurring at the output layer of the network. In contrast, the final layers of neural networks primarily function to cluster and associate these abstract patterns with the sample’s label.
@@ IV-E Analysis of Latent Space of Adversarially Robust Models
TABLE IX: Multi-category evaluation on samples transformed with image crop by using statistical metrics across object categories. Metrics include Mean ($\mu_{k^{*}}$), Standard deviation ($\sigma_{k^{*}}$), Skewness Coefficient ($\gamma_{k^{*}}$), Accuracy (Acc), and prevailing Pattern (Pat) in the context of diverse visual patterns.
@@
TABLE X: Performance on samples transformed with image crop across varied visual patterns - Pattern A (🟊normal-$\bigstar$), Pattern B (normal-$\clubsuit$), Pattern C (normal-$\spadesuit$), and $1,000$ ImageNet1k Classes. The table presents key statistics, including an average of Mean ($\mu_{k^{*}}$), Skewness coefficient ($\gamma_{k^{*}}$), Accuracy (Acc), and Number of Classes (N).
Average of Classes with Pattern A (🟊normal-$\bigstar$) (Fractured) Average of Classes with Pattern B (normal-$\clubsuit$) (Overlapped) Average of Classes with Pattern C (normal-$\spadesuit$) (Clustered) Average of 1,000 ImageNet1k Classes
@@
Having explored the changes in representation space with variations in training distribution, it is essential to address the susceptibility of models to adversarial attacks. To enhance robustness against such attacks, adversarially trained models have been proposed, incorporating adversarial samples in the training distribution [48]. In order to evaluate the shifts in the learned latent space between adversarially trained models and their non-robust counterparts, we examine the latent space of different robust models alongside their non-robust counterparts, as presented in Tables VII and VIII.
Observations from the table indicate that adversarially trained models tend to exhibit a more fractured distribution of samples in latent space compared to their non-robust counterparts. This explains the current tradeoff that exists between accuracy and robustness in the image classification models studied by Tsipras et al. [56]. This suggests that, in an effort to achieve clustering of heterogeneous feature samples within a single cluster as supervised with the pre-defined class label, models often compromise on learning robust features. Additionally, we observe a consistent trend where accuracy is associated with higher $\mu_{k^{*}}$ and lower $\gamma_{k^{*}}$, explaining why adversarially trained models may demonstrate reduced performance on clean samples compared to their non-robust counterparts.
IV-F Analysis of Latent Space of Different Input Transformations
TABLE XI: Multi-category evaluation on samples transformed with image rotation by using statistical metrics across object categories. Metrics include Mean ($\mu_{k^{*}}$), Standard deviation ($\sigma_{k^{*}}$), Skewness Coefficient ($\gamma_{k^{*}}$), Accuracy (Acc), and prevailing Pattern (Pat) in the context of diverse visual patterns.
@@ IV-F Analysis of Latent Space of Different Input Transformations
TABLE XII: Performance on samples transformed with image rotation across varied visual patterns - Pattern A (🟊normal-$\bigstar$), Pattern B (normal-$\clubsuit$), Pattern C (normal-$\spadesuit$), and $1,000$ ImageNet1k Classes. The table presents key statistics, including an average of Mean ($\mu_{k^{*}}$), Skewness coefficient ($\gamma_{k^{*}}$), Accuracy (Acc), and Number of Classes (N).
Average of Classes with Pattern A (🟊normal-$\bigstar$) (Fractured) Average of Classes with Pattern B (normal-$\clubsuit$) (Overlapped) Average of Classes with Pattern C (normal-$\spadesuit$) (Clustered) Average of 1,000 ImageNet1k Classes
@@
TABLE XIII: Multi-category evaluation on samples transformed with Gaussian noise by using statistical metrics across object categories. Metrics include Mean ($\mu_{k^{*}}$), Standard deviation ($\sigma_{k^{*}}$), Skewness Coefficient ($\gamma_{k^{*}}$), Accuracy (Acc), and prevailing Pattern (Pat) in the context of diverse visual patterns.
@@
TABLE XIV: Performance on samples transformed with Gaussian noise across varied visual patterns - Pattern A (🟊normal-$\bigstar$), Pattern B (normal-$\clubsuit$), Pattern C (normal-$\spadesuit$), and $1,000$ ImageNet1k Classes. The table presents key statistics, including an average of Mean ($\mu_{k^{*}}$), Skewness coefficient ($\gamma_{k^{*}}$), Accuracy (Acc), and Number of Classes (N).
Average of Classes with Pattern A (🟊normal-$\bigstar$) (Fractured) Average of Classes with Pattern B (normal-$\clubsuit$) (Overlapped) Average of Classes with Pattern C (normal-$\spadesuit$) (Clustered) Average of 1,000 ImageNet1k Classes
@@
TABLE XV: Multi-category evaluation of samples transformed with image stylization on models trained with different training distributions by using statistical metrics across object categories. Metrics include Mean ($\mu_{k^{*}}$), Standard deviation ($\sigma_{k^{*}}$), Skewness Coefficient ($\gamma_{k^{*}}$), Accuracy (Acc), and prevailing Pattern (Pat) in the context of diverse visual patterns.
@@
TABLE XVI: Performance by samples transformed with image stylization on models trained with different training distributions across varied visual patterns - Pattern A (🟊normal-$\bigstar$), Pattern B (normal-$\clubsuit$), Pattern C (normal-$\spadesuit$), and $1,000$ ImageNet1k Classes. The table presents key statistics, including an average of Mean ($\mu_{k^{*}}$), Skewness coefficient ($\gamma_{k^{*}}$), Accuracy (Acc), and Number of Classes (N).
Average of Classes with Pattern A (🟊normal-$\bigstar$) (Fractured) Average of Classes with Pattern B (normal-$\clubsuit$) (Overlapped) Average of Classes with Pattern C (normal-$\spadesuit$) (Clustered) Average of 1,000 ImageNet1k Classes
@@
TABLE XVII: Multi-category evaluation of adversarial samples created with PGD Attack by using statistical metrics across object categories. Metrics include Mean ($\mu_{k^{*}}$), Standard deviation ($\sigma_{k^{*}}$), Skewness Coefficient ($\gamma_{k^{*}}$), Accuracy (Acc), and prevailing Pattern (Pat) in the context of diverse visual patterns.
@@
TABLE XVIII: Performance on adversarial samples created with PGD Attack across varied visual patterns - Pattern A (🟊normal-$\bigstar$), Pattern B (normal-$\clubsuit$), Pattern C (normal-$\spadesuit$), and $1,000$ ImageNet1k Classes. The table presents key statistics, including an average of Mean ($\mu_{k^{*}}$), Skewness coefficient ($\gamma_{k^{*}}$), Accuracy (Acc), and Number of Classes (N).
Average of Classes with Pattern A (🟊normal-$\bigstar$) (Fractured) Average of Classes with Pattern B (normal-$\clubsuit$) (Overlapped) Average of Classes with Pattern C (normal-$\spadesuit$) (Clustered) Average of 1,000 ImageNet1k Classes
@@
The learned latent space of neural networks is known to be highly susceptible to slight perturbations in input samples, making it brittle. To comprehend the distribution of such transformed samples within the learned space, we analyze the latent space of ResNet-50 using samples modified with various input transformations. Specifically, we apply five different transformations, described below, to the input sample and individually evaluate their impact. Each sample is transformed once using the transformation to estimate their distribution in the latent space.
Image Crop:
In this transformation, the input sample is center-cropped to a size $s$. The reduced image is then resized to the standard $256\times 256$ dimensions and processed by the neural network. The results of this transformation are presented in Tables IX and X. Notably, as more pixels are cropped from the samples, the distribution of transformed samples becomes more fractured in the latent space. This suggests that the model relies on detecting multiple features within the image, and missing information can lead the model to perceive the sample as entirely different from the original.
Image Rotation:
[…]
Results of this transformation are presented in Tables XI and XII. Our examination of the distributions using k* distribution of rotated samples reveals a highly fractured distribution of transformed samples in the latent space implying that rotated samples are interpreted as distinct from non-rotated samples, also evident from the performance, as measured in accuracy.
Gaussian Noise
Image Stylization:
In this transformation, the input sample is deprived of its original texture and replaced with a random painting style, as illustrated by [53]. The results of this transformation on models trained on the standard ImageNet1k [45] dataset, a stylized version of ImageNet-1k (Stylized ImageNet) [53], and a combination of both are displayed in Tables XV and XVI.
Similar to other transformations, stylization in the images induces the fracturing of transformed samples. This further underscores the model’s sensitivity to variations in input samples, leading to distinct representations for stylized samples in the latent space.
Adversarial Perturbation:
[…]
The results of adversarial samples on robust and non-robust models are presented in Tables XVII and XVIII. Adversarial perturbations, like other transformations, cause fragmentation in the distribution of samples in the latent space. Although models trained with adversarial techniques exhibit increased robustness and improved transferability as demonstrated by Kotyan et al. [57], they are not completely robust, as indicated by a gradual rise in the fragmentation with stronger perturbations.
In summary, these findings highlight the neural network’s sensitivity to variations in input samples, showcasing that a model perceives transformed samples as substantially different from their original counterparts. The results across various input transformations, including adversarial perturbations, indicate that current models struggle to cluster transformed samples together, interpreting them as distinct not only from the original sample but also from each other.
IV-G Effect of Different Distance Metrics on k* Distribution
TABLE XIX: Multi-category evaluation using different distance metrics by using statistical metrics across object categories. Metrics include Mean ($\mu_{k^{*}}$), Standard deviation ($\sigma_{k^{*}}$), Skewness Coefficient ($\gamma_{k^{*}}$), and prevailing Pattern (Pat) in the context of diverse visual patterns.

### v2

@@ IV Experimental Results and Visualization
TABLE I: Multi-category evaluation of various neural architectures by using statistical metrics across object categories.
@@
TABLE II: Performance by various neural architectures across varied visual patterns.
Average of Classes with Pattern A (🟊$\bigstar$) (Fractured) Average of Classes with Pattern B ($\clubsuit$) (Overlapped) Average of Classes with Pattern C ($\spadesuit$) (Clustered) Average of 1,000 ImageNet1k Classes
@@
IV-A Common Experimental Setup
Datasets: Initially, we assess the k* distribution using the entire set of $1280$ images from the $16$-class-ImageNet dataset curated by Geirhos et al. [62]. This dataset contains $80$ images per each of the $16$ classes of the dataset. These are the $16$ entry-level categories from MS-COCO that have the highest number of ImageNet classes mapped via the WordNet hierarchy, making them compatible with the $1,000$ classes of ImageNet-1k dataset [63]. The purpose of employing this dataset is to evaluate individual classes organized naturally. In addition to the $16$-class-ImageNet dataset, we evaluate k* distribution comprehensively using $50,000$ validation images of the original ImageNet-1k dataset [63]. This dataset is divided into $1,000$ classes, each with $50$ images.
We apply various transformations to the samples from the $16$-class-ImageNet dataset and ImageNet-1k dataset to further scrutinize the models’ latent spaces. These transformations include Image Cropping, adding Gaussian Noise, rotating the images, and generating Adversarial and Stylized Versions of the samples. Note that the pre-trained models are not trained on these transformations.
Moreover, to assess the k* distribution across various other tasks, we also use various other datasets for different tasks. We use the English subset of the MASSIVE [64] dataset for intent classification. The testing samples contained $2,970$ samples split across $60$ distinct intent classes. For the keyword spotting task, we used the Speech Commands v0.02 dataset [65], which contains $4,890$ samples split across $36$ command categories. We utilize these datasets to analyze (see Figure 2) and comprehend the characteristics associated with each pattern (see Figure 1).
Deep Neural Network Architectures: We analyze the latent space of pre-trained weights from various open-source neural networks referenced as we have used them in the experiments.
Adversarial Attacks: We subject the trained ResNet-50 model to a Projected Gradient Descent (PGD) attack, as outlined in [66], employing a perturbation magnitude of ${\epsilon}=4/255$. This attack aims to gauge the extent of the effective alteration in representation induced by the adversarial samples. Furthermore, we assess the model that has undergone adversarial training using the PGD attack. This evaluation compares and contrasts the robust model’s latent space with its non-robust counterpart.
Metrics: For our analyses of individual classes, we report; Mean ($\mu_{k^{*}}$), Standard deviation ($\sigma_{k^{*}}$), Skewness Coefficient ($\gamma_{k^{*}}$), Accuracy (Acc), Number of Classes (N), and prevailing Pattern (Pat) in the context of diverse visual patterns.
TABLE III: Multi-category evaluation of various layers of ResNet-50 by using statistical metrics across object categories.
@@ IV-A Common Experimental Setup
TABLE IV: Performance by different layers of ResNet-50 across varied visual patterns.
Average of Classes with Pattern A (🟊$\bigstar$)  (Fractured) Average of Classes with Pattern B ($\clubsuit$)  (Overlapped) Average of Classes with Pattern C ($\spadesuit$)  (Clustered) Average of 1,000 ImageNet1k Classes
@@
IV-B Analysis of Latent Space of Different Neural Architectures
We have the option to visualize the learned latent space of ResNet-50 (logit layer) in two ways:
a) Using k* distribution, as illustrated in Figure 4 (Left), or,
b) By employing dimensionality reduction techniques, showcased in Figure 4 (Right).
Intriguingly, the t-SNE (Top Left) and UMAP (Bottom Right) visualizations indicate a highly clustered distribution of samples in latent space for ResNet-50, while Isomap (Top Right) and PCA (Bottom Left) show an overlapping distribution of samples in latent space. This creates uncertainty in distinguishing which classes are well-represented and which are fragmented in the distribution of samples in latent space.
To address this ambiguity and better understand the local latent space, we turn to the k* distribution, visualized in Figure 4 (Left). By utilizing the k* distribution, we can distinctly identify that the local spaces of the six visualized classes differ. For instance, the Airplane and Elephant classes exhibit more Pattern C ($\spadesuit$) clustered distribution of samples in latent space with a negatively skewed k* distribution.
In contrast, Bicycle and Dog classes showcase Pattern A (🟊$\bigstar$) fractured distribution of samples in latent space with a positively skewed k* distribution, while Car and Truck have Pattern B ($\clubsuit$) overlapped distribution of samples in latent space with an almost uniform k* distribution.
Furthermore, we can comprehensively compare various neural architectures, as presented in Table I. This allows us to observe how specific classes are distributed differently across architectures. Notably, the Bicycle class displays Pattern A (🟊$\bigstar$) (fractured distribution of samples) in ResNet and ResNeXt architectures, Pattern B ($\clubsuit$)  (overlapped distribution of samples) in ViT, and Pattern C ($\spadesuit$) (clustered distribution of samples) in EfficientNet-B0. Similarly, the Truck class is Pattern B ($\clubsuit$) (overlapped distribution of samples) in ResNet and ResNeXt architectures and Pattern C ($\spadesuit$) (clustered distribution of samples) in ViT and EfficientNet-B0. Conversely, Airplane, Dog, and Elephant classes exhibit consistent distribution across all architectures.
Additionally, a comprehensive comparison of neural architectures can be made by calculating the averages of Fractured, Overlapped, and Clustered classes, as detailed in Table II. The results indicate distinct distributions of classes across various architectures. For instance, ResNet tends to fracture the latent space, with only $9$ clustered classes. In contrast, ViT leans towards clustering of samples in the latent space, with a significant number of classes ($243$) exhibiting clustering, i.e., Pattern A (🟊$\bigstar$).
A general trend is also observed: an increase in the mean of k* distribution $\mu_{k^{*}}$ with improvements in model accuracy. Additionally, a declining general trend in the skewness coefficient of k* distribution $\gamma_{k^{*}}$ suggests a transition from fractured to the more overlapped distribution of samples in latent space across tested models.
IV-C Analysis of Latent Space of Different Layers of a Network
TABLE V: Multi-category evaluation of models trained on different distributions by using statistical metrics across object categories.
@@ IV-C Analysis of Latent Space of Different Layers of a Network
TABLE VI: Performance by models trained on different training datasets across varied visual patterns.
Average of Classes with Pattern A (🟊$\bigstar$) (Fractured) Average of Classes with Pattern B ($\clubsuit$) (Overlapped) Average of Classes with Pattern C ($\spadesuit$) (Clustered) Average of 1,000 ImageNet1k Classes
@@
TABLE VII: Multi-category evaluation of Robust (Adversarially Trained: AT) and Non-Robust (Standard Trained: ST) models by using statistical metrics across object categories.
@@
TABLE VIII: Performance of Robust (Adversarially Trained: AT) and Non-Robust (Standard Trained: ST) models across varied visual patterns.
Average of Classes with Pattern A (🟊$\bigstar$) (Fractured) Average of Classes with Pattern B ($\clubsuit$) (Overlapped) Average of Classes with Pattern C ($\spadesuit$) (Clustered) Average of 1,000 ImageNet1k Classes
@@
We analyze the latent space of various layers within ResNet-50 to gain insights into how classes are represented across different layers of the same model. Specifically, we examine the latent space of the final logit layer, the average pooling layer, and several convolution layers after different stages in the ResNet-50 architecture, as depicted in Tables III and IV.
Upon observation, we note that that number of overlapped classes from the average pooling layer decreases from $153$ to $56$ in the final logit layer, suggesting that logit layer fractures the overlapped regions. This is evident from the results as the number of fractured classes increases from $837$ in average pooling layer to $935$ for the logit layer.
@@ IV-E Analysis of Latent Space of Adversarially Robust Models
TABLE IX: Multi-category evaluation on samples transformed with image crop by using statistical metrics across object categories.
@@
TABLE X: Performance on samples transformed with image crop across varied visual patterns.
Average of Classes with Pattern A (🟊$\bigstar$) (Fractured) Average of Classes with Pattern B ($\clubsuit$) (Overlapped) Average of Classes with Pattern C ($\spadesuit$) (Clustered) Average of 1,000 ImageNet1k Classes
@@
TABLE XI: Multi-category evaluation on samples transformed with image rotation by using statistical metrics across object categories.
@@
Having explored the changes in representation space with variations in training distribution, it is essential to address models’ susceptibility to adversarial attacks. To enhance robustness against such attacks, adversarially trained models have been proposed, incorporating adversarial samples in the training distribution [66]. In order to evaluate the shifts in the learned latent space between adversarially trained models and their non-robust counterparts, we examine the latent space of different robust models alongside their non-robust counterparts, as presented in Tables VII and VIII.
Observations from the table indicate that adversarially trained models tend to exhibit a more fractured distribution of samples in latent space than their non-robust counterparts. This explains the current trade-off between accuracy and robustness in the image classification models studied by Tsipras et al. [70]. This suggests that, in an effort to achieve clustering of heterogeneous feature samples within a single cluster as supervised with the pre-defined class label, models often compromise on learning robust features. Additionally, we observe a consistent trend where accuracy is associated with higher $\mu_{k^{*}}$ and lower $\gamma_{k^{*}}$, explaining why adversarially trained models may demonstrate reduced performance on clean samples compared to their non-robust counterparts.
IV-F Analysis of Latent Space of Different Input Transformations
TABLE XII: Performance on samples transformed with image rotation across varied visual patterns.
Average of Classes with Pattern A (🟊$\bigstar$) (Fractured) Average of Classes with Pattern B ($\clubsuit$) (Overlapped) Average of Classes with Pattern C ($\spadesuit$) (Clustered) Average of 1,000 ImageNet1k Classes
@@ IV-F Analysis of Latent Space of Different Input Transformations
TABLE XIII: Multi-category evaluation on samples transformed with Gaussian noise by using statistical metrics across object categories.
@@
TABLE XIV: Performance on samples transformed with Gaussian noise across varied visual patterns.
Average of Classes with Pattern A (🟊$\bigstar$) (Fractured) Average of Classes with Pattern B ($\clubsuit$) (Overlapped) Average of Classes with Pattern C ($\spadesuit$) (Clustered) Average of 1,000 ImageNet1k Classes
@@
TABLE XV: Multi-category evaluation of samples transformed with image stylization on models trained with different distributions by using statistical metrics across object categories.
@@
TABLE XVI: Performance by samples transformed with image stylization on models trained with different distributions across varied visual patterns.
Average of Classes with Pattern A (🟊$\bigstar$) (Fractured) Average of Classes with Pattern B ($\clubsuit$) (Overlapped) Average of Classes with Pattern C ($\spadesuit$) (Clustered) Average of 1,000 ImageNet1k Classes
@@
The learned latent space of neural networks is known to be highly susceptible to slight perturbations in input samples, making it brittle. We analyze the latent space of ResNet-50 using samples modified with various input transformations to comprehend the distribution of such transformed samples. Specifically, we apply the five transformations described below to the input sample and individually evaluate their impact. Each sample is transformed once using the transformation to estimate their distribution in the latent space.
Image Crop:
The input sample is center-cropped to a size $s$ in this transformation. The reduced image is then resized to the standard $256\times 256$ dimensions and processed by the neural network. The results of this transformation are presented in Tables IX and X. Notably, as more pixels are cropped from the samples, the distribution of transformed samples becomes more fractured in the latent space. This suggests that the model relies on detecting multiple features within the image, and missing information can lead the model to perceive the sample as entirely different from the original.
Image Rotation:
[…]
Gaussian Noise
Image Stylization:
In this transformation, the input sample is deprived of its original texture and replaced with a random painting style [67]. The results of this transformation on models trained on the standard ImageNet-1k [63] dataset, a stylized version Stylized ImageNet [67], and a combination of both are displayed in Tables XV and XVI. Similar to other transformations, stylization in the images induces the fracturing of transformed samples. This further underscores the model’s sensitivity to variations in input samples, leading to distinct representations for stylized samples.
TABLE XVII: Multi-category evaluation of adversarial samples created with PGD Attack by using statistical metrics across object categories.
@@
TABLE XVIII: Performance on adversarial samples created with PGD Attack across varied visual patterns.
Average of Classes with Pattern A (🟊$\bigstar$) (Fractured) Average of Classes with Pattern B ($\clubsuit$) (Overlapped) Average of Classes with Pattern C ($\spadesuit$) (Clustered) Average of 1,000 ImageNet1k Classes
@@
TABLE XIX: Multi-category evaluation using different distance metrics by using statistical metrics across object categories.